OpenStudy (anonymous):

The sum of two numbers is one hundred sixty five. Their difference is sixty five. What are the numbers?

OpenStudy (anonymous):

do you know what sum and difference mean?

OpenStudy (anonymous):

@vvercetti

OpenStudy (anonymous):

@atreyu6s3x6 to add and subtract

OpenStudy (anonymous):

x+y=165, x-y=65 (115, 50)